OpenStudy (anonymous):

Please help! Medal will be given! Solve the equation for x. –10 + 3x + 6x = –10

OpenStudy (anonymous):

-10+9x=-10 +10 +10 9x=0 /9 X=0?

OpenStudy (anonymous):

I really don't get it :/ ...so after add 6 to each side how do you get the answer?

OpenStudy (anonymous):

Don't add 6, you add the 6x to the 3x giving you 9x

OpenStudy (anonymous):

Yeah, you're right @Toasterino

OpenStudy (anonymous):

Thank you! :)

OpenStudy (anonymous):

Np!